# ORCID: $0$-0002-6512-4229 (Om Prakash)

### v1

@@ Abstract
Let $\mathbb{F}_{q}$ be the finite field of $q=p^{m}$ elements where $p$ is a prime and $m$ is a positive integer. This paper considers $(\gamma,\Delta)$-cyclic codes over a class of finite commutative non-chain rings $\mathscr{R}_{q,s}=\mathbb{F}_{q}[v_{1},v_{2},\dots,v_{s}]/\langle v_{i}-v_{i}^{2},v_{i}v_{j}=v_{j}v_{i}=0\rangle$ where $\gamma$ is an automorphism of $\mathscr{R}_{q,s}$, $\Delta$ is a $\gamma$-derivation of $\mathscr{R}_{q,s}$ and $1\leq i\neq j\leq s$ for a positive integer $s$. Here, we show that a $(\gamma,\Delta)$-cyclic code of length $n$ over $\mathscr{R}_{q,s}$ is the direct sum of $(\theta,\Im)$-cyclic codes of length $n$ over $\mathbb{F}_{q}$, where $\theta$ is an automorphism of $\mathbb{F}_{q}$ and $\Im$ is a $\theta$-derivation of $\mathbb{F}_{q}$. Further, necessary and sufficient conditions for both $(\gamma,\Delta)$-cyclic and $(\theta,\Im)$-cyclic codes to contain their Euclidean duals are established. Finally, we obtain many quantum codes by applying the dual containing criterion on the Gray images of these codes. The obtained codes have better parameters than those available in the literature.
Keywords:
Skew polynomial rings Skew cyclic codes $(\sigma,\delta)$-cyclic codes Gray map CSS construction Quantum codes .
MSC:
12Y05 16Z05 94B05 94B35 94B15.
journal:
@@ 1 Introduction
After the pioneering work of Hammons et al. Hammons in 1994, codes over finite rings attracted many researchers for better error-correcting codes. Later, several important research has been carried out over finite rings and explored plenty of suitable parameters; we refer Alahmadi21 ; Ashraf19 ; Dertli15 ; Gao19 ; Ma18 . Nevertheless, all of these works have been considered over finite commutative rings. Hence, it is natural to look at these works over the noncommutative ring to obtain codes with better parameters. Towards this, in 2007, Boucher et al. Boucher1 introduced skew cyclic codes as a generalized class of cyclic codes by using a non-trivial automorphism $\theta$ on a finite field $\mathbb{F}_{q}$. They proved that the noncommutative rings (skew polynomial rings) are worthy alphabets for producing new parameters. In addition, they have provided a few codes with better parameters that are not known earlier over finite commutative rings. Since the factorization of the polynomial $x^{n}-1$ plays an important role in the characterization of cyclic codes of length $n$, more factorization leads to the case of getting many new code parameters. Therefore, skew cyclic codes that generalize cyclic codes in a noncommutative setup attract many researchers. Later, Abualrub et al. Abualrub and Bhaintwal Bhaintwal introduced skew-quasi cyclic codes. From an application point of view, recently, many authors have shown that skew cyclic codes are one of the important resources for producing new quantum codes along with classical codes (see Boucher1 ; Boucher11 ; Boucher2 ; Gursoy ; Jitman ; Siap ).
@@
However, all of the above works have been carried out on skew polynomial rings of automorphism type. Only a few works are available in the literature with both automorphisms and derivations. In Boucher ; MH ; P ; Luis , the authors generalized the notion of codes over skew polynomial rings with non-trivial automorphism $\theta$ and $\theta$- derivation $\Im$ under the usual addition of polynomials and a specific polynomials multiplication involving $\theta$ and $\Im$. For the noncommutative ring $\mathbb{F}_{q}[x,\theta;\Im]$ where $\theta$ is the Frobenius automorphism $a\mapsto a^{p}$, $p$ is the characteristic of $\mathbb{F}_{q}$, the authors Boucher ; Luis defined the inner $\theta$-derivation $\Im$ induced by $\beta\in\mathbb{F}_{q}^{*}$ of the form $a\mapsto\beta(\theta(a)-a)$. Further, Boulagouaz and Leroy MH studied $(\sigma,\delta)$-codes with $\sigma$-derivation induced by the ring element. Recently, Sharma and Bhaintwal Sharma have studied skew cyclic codes over $\mathbb{Z}_{4}+u\mathbb{Z}_{4},$ $u^{2}=1$ with both automorphism and inner derivation. In 2021, Ma et al. Li studied $(\sigma,\delta)$-skew quasi-cyclic codes over the ring $\mathbb{Z}_{4}+u\mathbb{Z}_{4},$ $u^{2}=1$. Further, in 2021, Patel and Prakash Patel21 studied $(\theta,\delta_{\theta})$-cyclic codes over the ring $\mathbb{F}_{q}[u,v]/\langle u^{2}-u,v^{2}-v,uv-vu\rangle$ via the decomposition method over $\mathbb{F}_{q}$. Here, we extend our previous work Patel21 to a more general structure and propose a fruitful application of $(\gamma,\Delta)$-cyclic codes in the context of quantum code construction. Per our survey, it is worth mentioning that this is the first article proposing an application of $(\gamma,\Delta)$-cyclic codes into quantum codes.
@@
Quantum error-correcting codes play a significant role in protecting information against the disturbances such as decoherence occurring in the channel. Shor Shor discovered the first quantum code. After that, Calderbank et al. Calderbank98 provided a method to obtain quantum codes from classical codes. This technique became very popular among researchers and is known as the CSS (Calderbank-Shor-Steane) construction. Presently, quantum codes and their implementation from classical codes have gained significant attention. As a consequence, many quantum codes with better parameters have been constructed from different families of linear codes such as cyclic, skew cyclic, skew constacyclic codes, etc., see Alahmadi21 ; Ashraf19 ; Dertli15 ; Gao19 ; Prakash21 ; Li20 ; Ma18 ; Ozen19 ; Prakash . But the searching of new methods on different structures are still going on by which one can construct quantum codes efficiently with suitable parameters. Since getting new quantum codes proportionally depend on the abundance of factors of $x^{n}-1$, many authors have been exploring quantum codes in the setting of the skew polynomial ring with automorphism where $x^{n}-1$ indeed possesses more factorization than the commutative case. However, in this work, we extend all these works in a new direction by considering skew polynomial rings with non-trivial automorphisms and derivations. In this case, we have different derivations for the same Frobenius automorphism having the form $a\mapsto\beta(\theta(a)-a)$ for all $\beta\in\mathbb{F}_{q}^{*}$.
@@ 2 Preliminaries
In this Section, we provide some preliminary results, definitions and notations which are used throughout this paper. We consider a finite non-chain ring $\mathscr{R}_{q,s}:=\mathbb{F}_{q}[v_{1},v_{2},\dots,v_{s}]/\langle v_{i}\\ -v_{i}^{2},v_{i}v_{j}=v_{j}v_{i}=0\rangle$ where $1\leq i\neq j\leq s$ and $s$ is a positive integer. This $\mathscr{R}_{q,s}$ is a class of finite commutative ring with unity for different values of $q$ and $s$. Further, $\mathscr{R}_{q,s}$ can also be represented in the form of $\mathscr{R}_{q,s}=\mathbb{F}_{q}+v_{1}\mathbb{F}_{q}+\cdots+v_{s}\mathbb{F}_{q}$ with $v_{i}-v_{i}^{2},v_{i}v_{j}=v_{j}v_{i}=0$. Moreover, $\mathscr{R}_{q,s}$ is a non-chain semi-local Frobenius ring having $s+1$ maximal ideals. For $s=2$, there are three maximal ideals $\langle v_{1}+v_{2}\rangle$, $\langle 1-v_{1}\rangle$ and $\langle 1-v_{2}\rangle$ in $\mathscr{R}_{q,2}$, refer Islam20b. Consider
$\zeta_{0}=\displaystyle\prod_{i=1}^{s}(1-v_{i}),~{}~{}\text{and}~{}~{}\zeta_{j}=v_{j},~{}~{}1\leq j\leq s.$
@@
$\zeta_{i}\zeta_{j}=\begin{cases}\zeta_{i},&\mbox{i}f~{}i=j\\ 0,&\mbox{i}f~{}i\neq j\end{cases}.$
Therefore, by Chinese Remainder Theorem, $\mathscr{R}_{q,s}=\zeta_{0}\mathscr{R}_{q,s}\oplus\zeta_{1}\mathscr{R}_{q,s}\oplus\cdots\oplus\zeta_{s}\mathscr{R}_{q,s}=\zeta_{0}\mathbb{F}_{q}\oplus\zeta_{1}\mathbb{F}_{q}\oplus\cdots\oplus\zeta_{s}\mathbb{F}_{q}$. From here we can conclude that any element $t\in\mathscr{R}_{q,s}$ can be uniquely written as $t=\zeta_{0}t_{0}+\zeta_{1}t_{1}+\cdots+\zeta_{s}t_{s}$, where $t_{i}\in\mathbb{F}_{q}$. Also, $t$ is a unit in $\mathscr{R}_{q,s}$ if and only if $t_{i}\in\mathbb{F}_{q}^{*}$ for all $i$.
Recall that a non-empty subset $\mathcal{C}$ of $\mathscr{R}_{q,s}^{n}$ is said to be a linear code of length $n$ over $\mathscr{R}_{q,s}$ if it is an $\mathscr{R}_{q,s}$-submodule of $\mathscr{R}_{q,s}^{n}$ and the elements of $\mathcal{C}$ are called codewords. The Hamming weight $w_{H}(c)$ of a codeword $c=(c_{0},c_{1},\dots,c_{n-1})\in\mathcal{C}$ is the number of nonzero coordinates in $c$. The Hamming distance between any two codewords $c$ and $c^{\prime}$ of $\mathcal{C}$ is defined as $d_{H}(c,c^{\prime})=w_{H}(c-c^{\prime})$ and the Hamming distance of a linear code $\mathcal{C}$ is defined as $d_{H}(\mathcal{C})=\min\{d_{H}(x,y)~{}|~{}x,~{}y\in\mathcal{C},x\neq y\}.$ The Euclidean inner product of $c$ and $c^{\prime}$ in $\mathcal{R}^{n}$ is defined by $c\cdot c^{\prime}=\sum_{i=0}^{n-1}c_{i}c^{\prime}_{i}$ where $c=(c_{0},c_{1},\dots,c_{n-1})$ and $c^{\prime}=(c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-1})$ are codewords in $\mathcal{C}$. The dual code of $\mathcal{C}$ is defined by $\mathcal{C}^{\perp}=\{c\in\mathscr{R}_{q,s}^{n}~{}|~{}c\cdot c^{\prime}=0,~{}\text{for all}~{}c^{\prime}\in\mathcal{C}\}$. Also, a linear code $\mathcal{C}$ is self-orthogonal if $\mathcal{C}\subseteq\mathcal{C}^{\perp}$ and self-dual if $\mathcal{C}=\mathcal{C}^{\perp}$. Further, let $c=(c_{0},c_{1},\dots,c_{n-1})\in\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$. If $\mathcal{C}$ is an $[n,k,d]$ linear code, then from the Singleton bound, its minimum distance is bounded above by $d\leq n-k+1$, where $d$ is the minimum distance, $k$ is the dimension, and $n$ is the length of the code. A code achieving the mentioned bound is called maximum-distance-separable (MDS). If the minimum distance of the code is one unit less than the MDS, then the code is called almost MDS. A linear code is said to be optimal if it has the highest possible minimum distance for a given length and dimension.
Definition 1
@@
$\mathscr{R}_{q,s}[x;\gamma,\Delta]=\{b_{l}x^{l}+\cdots+b_{1}x+b_{0}~{}|~{}b_{i}\in\mathscr{R}~{}\text{and}~{}l\in\mathbb{N}\},$
@@
Definition 3
N ; Leroy A pseudo-linear transformation $T_{\gamma,\Delta}:\mathscr{R}_{q,s}^{n}\rightarrow\mathscr{R}_{q,s}^{n}$ is an additive map defined by
@@
Remark 1
(KR, , Exercise 20) Let $\mathscr{R}_{q,s}[x;\gamma,\Delta]$ be a skew polynomial ring, $r\in\mathscr{R}_{q,s}$ and $n\in\mathbb{N}$. Then
@@
To find generator polynomials of $(\gamma,\Delta)$-cyclic codes over $\mathscr{R}_{q,s},$ first we derive the right division algorithm in $\mathscr{R}_{q,s}[x;\gamma,\Delta]$.
Theorem 2.1
( The Right Division Algorithm) Let $f(x),~{}g(x)\in\mathscr{R}_{q,s}[x;\gamma,\Delta]$ such that the leading coefficient of $g(x)$ be a unit. Then there exist $q(x),~{}r(x)\in\mathscr{R}_{q,s}[x;\gamma,\Delta]$ such that
@@
where $r(x)=0$ or $\deg~{}r(x)<\deg~{}g(x)$.
Proof
If $f(x)=0$, then result follows by taking $q(x),~{}r(x)=0$. If $\deg~{}f(x)<\deg~{}g(x)$, then we take $q(x)=0$ and $r(x)=f(x)$. Further, for $\deg f(x)\geq\deg g(x)$, we prove it by induction on $\deg f(x)$.
It can be easily seen that the result is true for $\deg f(x)=0$. Now, suppose the result is true for all polynomials of degree less than $\deg~{}f(x)$. Let $f(x)=f_{0}+f_{1}x+\cdots+f_{s}x^{s}$ and $g(x)=g_{0}+g_{1}x+\cdots+g_{t}x^{t}$ be two polynomials in $\mathscr{R}_{q,s}[x;\gamma,\Delta]$ such that $f_{s}\neq 0$ and $g_{t}$ is a unit. Consider a polynomial
@@
From Remark 1, $h(x)$ can be written as
@@
where $a_{1},a_{2},\dots,a_{s-t-1},b_{1},b_{2},\dots,b_{s-t-1}\in\mathscr{R}_{q,s}$. Now, we can conclude that $\deg~{}h(x)<\deg~{}f(x)$. Hence, by induction on $\deg~{}h(x),$ there exist $b(x),r(x)\in\mathscr{R}_{q,s}[x;\gamma,\Delta]$ such that
@@
where $r(x)=0$ or $\deg~{}r(x)<\deg~{}g(x)$. Thus,
@@
where $q(x)=b(x)+f_{s}\gamma^{s-t}(g_{t}^{-1})x^{s-t}\in\mathscr{R}_{q,s}[x;\gamma,\Delta]$ and $r(x)=0$ or $\deg~{}r(x)<\deg~{}g(x)$. This gives the required result.

### v2

# Encoding and Construction of Quantum Codes from $(\gamma,\Delta)$-cyclic Codes over a Class of Non-chain Rings

@@ Abstract
Let $\mathbb{F}_{q}$ be a finite field of $q=p^{m}$ elements where $p$ is a prime and $m$ is a positive integer. This paper considers $(\gamma,\Delta)$-cyclic codes over a class of finite non-chain commutative rings $\mathscr{R}_{q,s}=\mathbb{F}_{q}[v_{1},v_{2},\dots,v_{s}]/\langle v_{i}-v_{i}^{2},v_{i}v_{j}=v_{j}v_{i}=0\rangle$ where $\gamma$ is an automorphism of $\mathscr{R}_{q,s}$, $\Delta$ is a $\gamma$-derivation of $\mathscr{R}_{q,s}$ and $1\leq i\neq j\leq s$ for a positive integer $s$. Here, we show that a $(\gamma,\Delta)$-cyclic code of length $n$ over $\mathscr{R}_{q,s}$ is the direct sum of $(\theta,\Im)$-cyclic codes of length $n$ over $\mathbb{F}_{q}$, where $\theta$ is an automorphism of $\mathbb{F}_{q}$ and $\Im$ is a $\theta$-derivation of $\mathbb{F}_{q}$. Further, necessary and sufficient conditions for both $(\gamma,\Delta)$-cyclic and $(\theta,\Im)$-cyclic codes to contain their Euclidean duals are established. Then, we obtain many quantum codes by applying the dual containing criterion on the Gray images of these codes. These codes have better parameters than those available in the literature. Finally, the encoding and error-correction procedures for our proposed quantum codes are discussed.
Keywords: Skew polynomial rings, skew cyclic codes, $(\sigma,\delta)$-cyclic codes, Gray map,
CSS construction, quantum codes.
MSC (2020): 12L10 $\cdot$ 16Z05 $\cdot$ 94B05 $\cdot$ 94B35 $\cdot$ 94B15.
@@ 1 Introduction
After the pioneering work of Hammons et al. [22] in 1994, codes over finite rings attracted many researchers for better error-correcting codes. Later, several important research has been carried out over finite rings and explored plenty of suitable parameters; we refer [2, 3, 13, 16, 33]. Nevertheless, all of these works have been considered over finite commutative rings. Hence, it is natural to look at these works over the noncommutative ring to obtain codes with better parameters. Towards this, in 2007, Boucher et al. [5] introduced skew cyclic codes as a generalized class of cyclic codes using a non-trivial automorphism $\theta$ on a finite field $\mathbb{F}_{q}$. They proved that the noncommutative rings (skew polynomial rings) are worthy alphabets for producing new parameters. In addition, they have provided a few codes with better parameters that were not known earlier over finite commutative rings. The factorization of the polynomial $x^{n}-1$ plays an important role in the characterization of cyclic codes of length $n$ and more factorization leads to the case of getting many new codes with better parameters. Therefore, skew cyclic codes that generalize cyclic codes in a noncommutative setup attract many researchers. During 2010-2012, Abualrub et al. [1] and Bhaintwal [4] introduced and developed some interesting results on skew-quasi cyclic codes. From an application point of view, recently, many authors have shown that skew cyclic codes are one of the important resources for producing new quantum codes along with classical codes [5, 6, 8, 21, 26, 43].
@@
However, all of the above works have been carried out on skew polynomial rings of automorphism type. Only a few works are available in the literature with both automorphisms and derivations. In [7, 9, 39, 45], the authors generalized the notion of codes over skew polynomial rings with non-trivial automorphism $\theta$ and $\theta$- derivation $\Im$ under the usual addition of polynomials and a specific polynomials multiplication involving $\theta$ and $\Im$. For the noncommutative ring $\mathbb{F}_{q}[x,\theta;\Im]$ where $\theta$ is the Frobenius automorphism $a\mapsto a^{p}$, $p$ is the characteristic of $\mathbb{F}_{q}$, the authors [7, 45] defined the inner $\theta$-derivation $\Im$ induced by $\beta\in\mathbb{F}_{q}^{*}$ of the form $a\mapsto\beta(\theta(a)-a)$. Further, Boulagouaz and Leroy [9] studied $(\sigma,\delta)$-codes with $\sigma$-derivation induced by the ring element. Recently, Sharma and Bhaintwal [41] have studied skew cyclic codes over $\mathbb{Z}_{4}+u\mathbb{Z}_{4},$ $u^{2}=1$ with both automorphism and inner derivation. In 2021, Ma et al. [32] studied $(\sigma,\delta)$-skew quasi-cyclic codes over the ring $\mathbb{Z}_{4}+u\mathbb{Z}_{4},$ $u^{2}=1$. Further, in 2021, Patel and Prakash [38] studied $(\theta,\delta_{\theta})$-cyclic codes over the ring $\mathbb{F}_{q}[u,v]/\langle u^{2}-u,v^{2}-v,uv-vu\rangle$ via the decomposition method over $\mathbb{F}_{q}$. Here, we extend our previous work [38] to a more general structure and propose a fruitful application of $(\gamma,\Delta)$-cyclic codes in the context of quantum code construction. As per our survey, it is worth mentioning that this is the first article proposing an application of $(\gamma,\Delta)$-cyclic codes into quantum codes.
@@
Quantum error-correcting codes play a significant role in protecting information against disturbances such as decoherence occurring in the channel. In this connection, in 1995, Shor [42] discovered the first quantum code. After that, Calderbank et al. [11] provided a method to obtain quantum codes from classical codes. This technique became very popular among researchers and is known as the CSS (Calderbank-Shor-Steane) construction. Presently, quantum codes and their implementation from classical codes have gained significant attention. As a consequence, many quantum codes with better parameters have been constructed from different families of linear codes such as cyclic, skew cyclic, skew constacyclic codes, etc., see [2, 3, 13, 16, 30, 33, 37, 40]. However, the search for new methods on different structures is still ongoing by which one can construct quantum codes efficiently with suitable parameters. Since getting new quantum codes proportionally depends on the abundance of factors of $x^{n}-1$, many authors have been exploring quantum codes in the setting of the skew polynomial ring with automorphism where $x^{n}-1$ indeed possesses more factorization than the commutative case. Thus, in this work, we extend all these previous works in a new direction by considering skew polynomial rings with non-trivial automorphisms and nonzero derivations. Here, we use different derivations for the same Frobenius automorphism having the form $a\mapsto\beta(\theta(a)-a)$ for all $\beta\in\mathbb{F}_{q}^{*}$.
@@ 2 Preliminaries
In this Section, we provide some preliminary results, definitions and notations which are used throughout this paper. We consider a finite non-chain ring $\mathscr{R}_{q,s}:=\mathbb{F}_{q}[v_{1},v_{2},\dots,v_{s}]/\langle v_{i}-v_{i}^{2},v_{i}v_{j}=v_{j}v_{i}=0\rangle$ where $1\leq i\neq j\leq s$ and $s$ is a positive integer. This $\mathscr{R}_{q,s}$ is a class of finite commutative ring with unity for different values of $q$ and $s$. Further, $\mathscr{R}_{q,s}$ can also be represented in the form of $\mathscr{R}_{q,s}=\mathbb{F}_{q}+v_{1}\mathbb{F}_{q}+\cdots+v_{s}\mathbb{F}_{q}$ with $v_{i}-v_{i}^{2},v_{i}v_{j}=v_{j}v_{i}=0$. Moreover, $\mathscr{R}_{q,s}$ is a non-chain semi-local Frobenius ring having $s+1$ maximal ideals. For $s=2$, there are three maximal ideals $\langle v_{1}+v_{2}\rangle$, $\langle 1-v_{1}\rangle$ and $\langle 1-v_{2}\rangle$ in $\mathscr{R}_{q,2}$, refer [23]. Consider
$\zeta_{0}=\displaystyle\prod_{i=1}^{s}(1-v_{i}),\leavevmode\nobreak\ \leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ \leavevmode\nobreak\ \zeta_{j}=v_{j},\leavevmode\nobreak\ \leavevmode\nobreak\ 1\leq j\leq s.$
@@
$\zeta_{i}\zeta_{j}=\begin{cases}\zeta_{i},&\mbox{i}f\leavevmode\nobreak\ i=j\\ 0,&\mbox{i}f\leavevmode\nobreak\ i\neq j\end{cases}.$
Hence, by Chinese Remainder Theorem, $\mathscr{R}_{q,s}=\zeta_{0}\mathscr{R}_{q,s}\oplus\zeta_{1}\mathscr{R}_{q,s}\oplus\cdots\oplus\zeta_{s}\mathscr{R}_{q,s}=\zeta_{0}\mathbb{F}_{q}\oplus\zeta_{1}\mathbb{F}_{q}\oplus\cdots\oplus\zeta_{s}\mathbb{F}_{q}$. Thus, we conclude that any element $t\in\mathscr{R}_{q,s}$ can be uniquely written as $t=\zeta_{0}t_{0}+\zeta_{1}t_{1}+\cdots+\zeta_{s}t_{s}$, where $t_{i}\in\mathbb{F}_{q}$. Also, $t$ is a unit in $\mathscr{R}_{q,s}$ if and only if $t_{i}\in\mathbb{F}_{q}^{*}$ for all $i$.
Recall that a non-empty subset $\mathcal{C}$ of $\mathscr{R}_{q,s}^{n}$ is said to be a linear code of length $n$ over $\mathscr{R}_{q,s}$ if it is an $\mathscr{R}_{q,s}$-submodule of $\mathscr{R}_{q,s}^{n}$ and the elements of $\mathcal{C}$ are called codewords. The Hamming weight $w_{H}(c)$ of a codeword $c=(c_{0},c_{1},\dots,c_{n-1})\in\mathcal{C}$ is the number of nonzero coordinates in $c$. The Hamming distance between any two codewords $c$ and $c^{\prime}$ of $\mathcal{C}$ is defined as $d_{H}(c,c^{\prime})=w_{H}(c-c^{\prime})$ and the Hamming distance of a linear code $\mathcal{C}$ is defined as $d_{H}(\mathcal{C})=\min\{d_{H}(x,y)\leavevmode\nobreak\ |\leavevmode\nobreak\ x,\leavevmode\nobreak\ y\in\mathcal{C},x\neq y\}.$ The Euclidean inner product of $c$ and $c^{\prime}$ in $\mathcal{R}^{n}$ is defined by $c\cdot c^{\prime}=\sum_{i=0}^{n-1}c_{i}c^{\prime}_{i}$ where $c=(c_{0},c_{1},\dots,c_{n-1})$ and $c^{\prime}=(c^{\prime}_{0},c^{\prime}_{1},\dots,c^{\prime}_{n-1})$ are codewords in $\mathcal{C}$. The dual code of $\mathcal{C}$ is defined by $\mathcal{C}^{\perp}=\{c\in\mathscr{R}_{q,s}^{n}\leavevmode\nobreak\ |\leavevmode\nobreak\ c\cdot c^{\prime}=0,\leavevmode\nobreak\ \text{for all}\leavevmode\nobreak\ c^{\prime}\in\mathcal{C}\}$. Also, a linear code $\mathcal{C}$ is self-orthogonal if $\mathcal{C}\subseteq\mathcal{C}^{\perp}$ and self-dual if $\mathcal{C}=\mathcal{C}^{\perp}$. Further, let $c=(c_{0},c_{1},\dots,c_{n-1})\in\mathcal{C}\subseteq\mathbb{F}_{q}^{n}$. If $\mathcal{C}$ is an $[n,k,d]$ linear code, then from the Singleton bound, its minimum distance is bounded above by $d\leq n-k+1$, where $d$ is the minimum distance, $k$ is the dimension, and $n$ is the length of the code. A code achieving the mentioned bound is called maximum-distance-separable (MDS). If the minimum distance of the code is one unit less than the MDS, then the code is called almost MDS. A linear code is said to be optimal if it has the highest possible minimum distance for a given length and dimension.
Definition 2.1.
@@
$\mathscr{R}_{q,s}[x;\gamma,\Delta]=\{b_{l}x^{l}+\cdots+b_{1}x+b_{0}\leavevmode\nobreak\ |\leavevmode\nobreak\ b_{i}\in\mathscr{R}\leavevmode\nobreak\ \text{and}\leavevmode\nobreak\ l\in\mathbb{N}\},$
@@
Definition 2.3.
[25, 29] A pseudo-linear transformation $T_{\gamma,\Delta}:\mathscr{R}_{q,s}^{n}\rightarrow\mathscr{R}_{q,s}^{n}$ is an additive map defined by
@@
Remark 2.1.
[17, Exercise 20] Let $\mathscr{R}_{q,s}[x;\gamma,\Delta]$ be a skew polynomial ring, $r\in\mathscr{R}_{q,s}$ and $n\in\mathbb{N}$. Then
@@
To find generator polynomials of $(\gamma,\Delta)$-cyclic codes over $\mathscr{R}_{q,s},$, first we derive the right division algorithm in $\mathscr{R}_{q,s}[x;\gamma,\Delta]$.
Theorem 2.1.
( The Right Division Algorithm) Let $f(x),\leavevmode\nobreak\ g(x)\in\mathscr{R}_{q,s}[x;\gamma,\Delta]$ such that the leading coefficient of $g(x)$ be a unit. Then there exist $q(x),\leavevmode\nobreak\ r(x)\in\mathscr{R}_{q,s}[x;\gamma,\Delta]$ such that
@@
where $r(x)=0$ or $\deg\leavevmode\nobreak\ r(x)<\deg\leavevmode\nobreak\ g(x)$.
Proof.
If $f(x)=0$, then the result follows by taking $q(x),\leavevmode\nobreak\ r(x)=0$. If $\deg\leavevmode\nobreak\ f(x)<\deg\leavevmode\nobreak\ g(x)$, then we take $q(x)=0$ and $r(x)=f(x)$. Furthermore, for $\deg f(x)\geq\deg g(x)$, we prove it by induction on $\deg f(x)$.
It can be easily seen that the result is true for $\deg f(x)=0$. Now, suppose the result is true for all polynomials of degree less than $\deg\leavevmode\nobreak\ f(x)$. Let $f(x)=f_{0}+f_{1}x+\cdots+f_{s}x^{s}$ and $g(x)=g_{0}+g_{1}x+\cdots+g_{t}x^{t}$ be two polynomials in $\mathscr{R}_{q,s}[x;\gamma,\Delta]$ such that $f_{s}\neq 0$ and $g_{t}$ is a unit. Consider a polynomial
@@
From Remark 2.1, $h(x)$ can be written as
@@
where $a_{1},a_{2},\dots,a_{s-t-1},b_{1},b_{2},\dots,b_{s-t-1}\in\mathscr{R}_{q,s}$. Now, we can conclude that $\deg\leavevmode\nobreak\ h(x)<\deg\leavevmode\nobreak\ f(x)$. Hence, by induction on $\deg\leavevmode\nobreak\ h(x),$ there exist $b(x),r(x)\in\mathscr{R}_{q,s}[x;\gamma,\Delta]$ such that
@@
where $r(x)=0$ or $\deg\leavevmode\nobreak\ r(x)<\deg\leavevmode\nobreak\ g(x)$. Thus,
@@
where $q(x)=b(x)+f_{s}\gamma^{s-t}(g_{t}^{-1})x^{s-t}\in\mathscr{R}_{q,s}[x;\gamma,\Delta]$ and $r(x)=0$ or $\deg\leavevmode\nobreak\ r(x)<\deg\leavevmode\nobreak\ g(x)$. This gives the required result. ∎